solve for x: 1+3x/(x+5)(x-2)=2/(x+5)+3/(x-2)
Answer by olayinka003(5)

if i am on the line,this is your question:
1+3x/(x+5)(x-2)=2/(x+5)+3/(x-2).its a little bit long.anyway this is all i can offer to it:
1+3x/(x+5)(x-2)= 1+3x/(x^2 -2x + 5x -10)
2/(x+5)+3/(x-2)= 2(x-2) + 3(x+5) / x^2 -2x + 5x -10
so we therefore have:
from the left hand side: 1 + 3x / x^2 - 3x - 10.
then from the right hand side: 2x - 4 + 3x + 15 / x^2 3x -10
so the right hand side becomes: 5x + 11 / x^2 3x -10
the whole expression becomes:
1 + 3x / x^2 - 3x - 10 = 2x - 4 + 3x + 15 / x^2 - 3x -10
hope you are cathching it.
so lets go on.we have to cross multiply here.then we have
1 + 3x(x^2- 3x -10) = 5x + 11(x^2 - 3x -10)
so:you multiply through the bracket:
x^2 - 3x -10 + 3x^3 - 9x^2 - 30x = 5x^3 - 15x^2 - 50x +11x^2 - 33x - 110
x^2 - 3x -10 + 3x^3 - 9x^2 - 30x = 5x^3 - 3x^2 - 50x - 33x -110
so crossing the equality sign:all + and - on the right hand side side will change to thier opposites respectively:
we have:
x^2 - 33x -10 + 3x^3 - 9x^2 - 5x^3 + 3x^2 + 83x + 110
so x= 2x^2 + 100 + 50x - 2x^3
= 2x^3 + 2x^2 + 50x + 100
just take your time and go through it.u ll get the whole idea.